# Exploring the Roles of Large Language Models in Reshaping Transportation Systems: A Survey, Framework, and Roadmap

### v1

@@ 7.3 Spatial-temporal predictor
For macroscopic mobility prediction, LLMs are applied to decode complex spatial-temporal dependencies, contextual cues, and behavioral trends embedded in diverse datasets such as GPS trajectories, transit schedules, social media activity, and traffic reports (Liu et al., 2024g). Unlike traditional machine learning approaches, LLMs excel at synthesizing unstructured text with structured mobility data based on their fundamental knowledge, allowing them to model nuanced interactions between infrastructure, environment, and user behavior. To help LLMs analyze human mobility data, LLM-Mob introduced by Wang et al. (2023d) presents concepts of historical stays and context stays to capture both long-term and short-term dependencies in human movement. Additionally, context-inclusive prompts are designed to improve the accuracy of LLMs in generating time-aware predictions. Similarly, prompt-based prediction is also explored in LLM-MPE (Liang et al., 2024a). LLM-MPE pompts LLMs to process textual descriptions of public events and historical mobility data to predict human mobility during such events. It converts raw, unstructured event descriptions into a standardized format and segments historical mobility data into regular and event-related components in the prompt. Training foundation models using pure mobility data also demonstrates great potentials. In (Haydari et al., 2024), MobilityGPT is introduced as a geospatially-aware generative model that formulates human mobility modeling as an autoregressive generation task using the GPT architecture. They fine-tune MobilityGPT using a Reinforcement Learning from Trajectory Feedback (RLTF) mechanism, minimizing the travel distance between training and synthetically generated trajectories. UniMob (Long et al., 2024b) extents this paradigm and endeavors to unify individual trajectory and crowd flow predictions. UniMob employs a multi-view mobility tokenizer to transform both trajectory and flow data into spatiotemporal tokens, facilitating unified sequential modeling through a diffusion Transformer architecture. Finally, LLMs are also integrated into multimodal demand prediction by fusing heterogeneous data sources and reformulating time series forecasting, such as electric vehicles charging demand (Qu et al., 2024), taxi and bike usage demand (Liu et al., 2024b), and package delivery demand (Nie et al., 2025a).
@@ 7.4 Summary and outlook
The integration of LLMs as decision facilitators has introduced remarkable capabilities in transportation systems, spanning decision-making, action guidance, and spatial-temporal forecasting. As decision makers, LLMs demonstrate human-like adaptability in traffic signal control, autonomous driving, and route optimization, outperforming traditional methods while offering explainable reasoning. In guiding decisions, LLMs bridge high-level reasoning with low-level control systems through interpretable instructions, enhancing safety and efficiency in complex scenarios. For spatial-temporal forecasting, LLMs decode intricate mobility patterns by unifying multimodal data representations, achieving competitive performance in traffic flow and trajectory prediction through tokenization and fine-tuning strategies. As LLMs evolve from auxiliary tools to central predictors, their integration with ITS techniques such as digital twin platforms and IoT ecosystems will likely promote predictive intelligence in smart transportation applications.
@@
    Data representation and computation: LLMs’ text-centric training limits their effectiveness in processing numerical and geometric data inherent to transportation systems. Although tokenization methods such as spatial-temporal graph tokenizers (Liu et al., 2024d) and vector captioning (Chen et al., 2024b) show promise, fundamental gaps persist in representing continuous physical spaces and performing precise numerical computations. It remains an open question how to develop a customized data representation method that is suitable for multiscale traffic data (Nie et al., 2024b) and adaptable to LLMs.
@@
    Safety and alignment: Critical applications like autonomous driving require rigorous safety guarantees. Current frameworks address this through simulation sandboxes (Fu et al., 2024a) and safety alignment techniques (Yang et al., 2024c; Xie et al., 2025), but real-world deployment demands formal verification methods and enhanced robustness against adversarial inputs (Liu et al., 2024c). Equipping LLMs with safety evaluation modules is a crucial step before practical implementation.
@@ 8.2 Available resources
A Survey of LLM Surveys: https://github.com/NiuTrans/ABigSurveyOfLLMs 2024 GitHub Compiles 150+ surveys across subfields like alignment, safety, and multimodal LLMs.
@@ 8.3 Computational requirement
Pretraining a foundational LLM in vertical domain such as transportation can be infeasible due to extremely high resource consumption. However, recent advances in parameter-efficient fine-tuning methods such as LoRA (Hu et al., 2022) and QLoRA (Dettmers et al., 2023) have made customization of LLMs widely accessible, allowing mid-range hardware to handle models previously restricted to enterprise-grade infrastructure.
@@ 9.1 Future opportunities of LLM4TR
    Eliciting latent problem-solving abilities for emergent challenges: After being pre-trained on large-scale corpora, LLMs are possessed with potential abilities as general-purpose task solvers. These abilities might not be explicitly exhibited when LLMs perform some specific traffic problems. Therefore, it is useful to design suitable task instructions or specific in-context learning strategies to elicit such underexplored abilities (Shin et al., 2020; Wei et al., 2022b). Strategic ability elicitation could unlock these capacities. For example, role-playing or persona adoption strategy instructs the model to adopt the persona of a domain expert (e.g., an experienced traveler (Liu et al., 2024e)). This approach leverages the latent knowledge by framing responses in the voice and reasoning style of an expert, which can yield more domain-appropriate answers. The problem decomposition strategy breaks a complex domain-specific problem into a series of simpler sub-problems. Then the compositional tasks can be solved by meta agents. Finally, using domain-contextualized instruction tuning to guide LLMs to focus on relevant domain aspects and follow a structured approach to the problem.
@@
    Model compression for real-time decision-making: Deploying LLMs on resource-constrained edge devices (e.g., onboard vehicle computers, roadside units) necessitates lightweight yet capable models (Liu and Zhao, 2024). Recent advances in model compression suggests a promising pathway. By dynamically pruning non-critical parameters, e.g., tailoring sparsification to preserve traffic-related knowledge such as route optimization while compressing unrelated linguistic knowledge, models can be significantly reduced in size without undermining task-specific performance (Ma et al., 2023a). Furthermore, hardware-aware quantization methods (Lin et al., 2024a), such as mixed-precision and adaptive quantization schemes, are emerging to exploit the computational strengths of modern GPUs/TPUs. They can reduce memory footprint and inference latency in a way that is sensitive to the unique requirements of transportation applications, without reliance on cloud APIs.
@@ 9.2 Challenges and concerns
Despite the impressive advances in LLMs and their potential to transform transportation systems, the integration and deployment of LLMs into the safety-critical, ethically sensitive and computationally constrained domains raises several challenges that need careful attention. Below, we discuss these concerns related to model biases, domain limitations, computational requirements, data privacy, and ethical implications in depth.
@@
    Bias and hallucinations inherent in LLMs: LLMs are typically trained on massive, heterogeneous datasets collected from the internet. Consequently, they may inherit and even amplify biases present in the training data. In transportation contexts, such biases can lead to erroneous predictions or inappropriate recommendations. In addition, the presence of hallucinations, which can cause LLMs to generate factually incorrect information, poses significant risks. In safety-critical systems like autonomous vehicles or traffic control centers, even minor hallucinations can have severe real-world consequences.
@@
    Ensuring rigor and controllability in LLM-driven solutions: While LLMs offer great potential, their integration into transportation systems requires rigor in performance guarantees. A core challenge lies in reconciling the probabilistic, opaque nature of LLMs with the deterministic nature of traffic rules and reliability required for real-time traffic management. Researchers need to develop certifiable validation frameworks that quantify uncertainty bounds, enforce context consistency in model outputs, and rigorously test LLM-based solutions against adversarial scenarios. To maintain human supervision, explainability tools like counterfactual reasoning interfaces should be mandated, allowing operators to monitor how LLM-derived recommendations align with controllability and operational constraints.
@@
    Domain-specific problem-solving and numeric computing limitations: While LLMs excel at processing and generating natural language, their capabilities in precise numerical computing and domain-specific problem solving remain limited. Transportation problems often include precise and quantitative analysis that depend on accurate numerical computations or mathematical optimization, such as real-time traffic prediction, equilibrium analysis, and dynamic routing. Standard LLMs are not inherently designed for these tasks, which can lead to suboptimal performance when used in isolation.
@@
    Computational demand for adapting LLMs: The impressive performance of LLMs comes at a high computational cost. Even fine-tuning and adapting these models on user-defined datasets necessitates prohibitive computational resources. Previous data-driven transportation modeling typically required only moderate or small computational overhead. Thus, computation in the era of LLMs becomes a technical bottleneck for transportation researchers and practitioners with limited hardware computing capacity. In addition, while API-based solutions offer an accessible approach to using pre-trained LLMs provided by third-part companies, the limited access and model agnosticism largely limits the ways in which capabilities of LLMs can be explored and applied.
@@
    Privacy and security concerns: LLMs require large amounts of data, some of which may be sensitive or personal such as real-time location data, travel patterns, or driver behavior information. This raises significant privacy concerns when LLMs memorize and leak personal information. In addition to privacy issues, there are also security risks related with deploying LLMs as decision makers in critical infrastructure. For instance, cyberattacks using adversarial prompts can manipulate LLM-based traffic controller. Such adversarial vulnerabilities of LLMs are adverse for user’s security and system’s functionality.
@@
    Ethical, legal, and equitable concerns: Finally, the integration of LLMs in transportation systems raises several ethical, legal, and equitable issues. The black-box nature of many LLMs complicates interpretability, making it difficult to trace decisions or explain errors when they occur. This opacity is problematic in scenarios where transparency is required for public safety and trust, such as autonomous vehicles. Furthermore, there are concerns about equitable access. Biased outputs of LLMs may disproportionately affect underrepresented or vulnerable groups. For example, route optimization models may prioritize routes through lower-income neighborhoods as "acceptable" congestion zones. While the high cost of deploying and maintaining advanced LLMs may also widen the gap between well-funded urban centers and less affluent regions. Lastly, legal liability in the event of system failures remains largely undefined. If an LLM-recommended merging strategy causes a collision of autonomous vehicles, is the responsibility with the model developer, traffic agency, or vehicle original equipment manufacturer?

### v2

@@ 7.3 Spatial-temporal predictor
For macroscopic mobility prediction, LLMs are applied to decode complex spatial-temporal dependencies, contextual cues, and behavioral trends embedded in diverse datasets such as GPS trajectories, transit schedules, social media activity, and traffic reports (Liu et al., 2024h). Unlike traditional machine learning approaches, LLMs excel at synthesizing unstructured text with structured mobility data based on their fundamental knowledge, allowing them to model nuanced interactions between infrastructure, environment, and user behavior. To help LLMs analyze human mobility data, LLM-Mob introduced by Wang et al. (2023d) presents concepts of historical stays and context stays to capture both long-term and short-term dependencies in human movement. Additionally, context-inclusive prompts are designed to improve the accuracy of LLMs in generating time-aware predictions. Similarly, prompt-based prediction is also explored in LLM-MPE (Liang et al., 2024a). LLM-MPE pompts LLMs to process textual descriptions of public events and historical mobility data to predict human mobility during such events. It converts raw, unstructured event descriptions into a standardized format and segments historical mobility data into regular and event-related components in the prompt. Training foundation models using pure mobility data also demonstrates great potentials. In (Haydari et al., 2024), MobilityGPT is introduced as a geospatially-aware generative model that formulates human mobility modeling as an autoregressive generation task using the GPT architecture. They fine-tune MobilityGPT using a Reinforcement Learning from Trajectory Feedback (RLTF) mechanism, minimizing the travel distance between training and synthetically generated trajectories. UniMob (Long et al., 2024b) extents this paradigm and endeavors to unify individual trajectory and crowd flow predictions. UniMob employs a multi-view mobility tokenizer to transform both trajectory and flow data into spatiotemporal tokens, facilitating unified sequential modeling through a diffusion Transformer architecture. Finally, LLMs are also integrated into multimodal demand prediction by fusing heterogeneous data sources and reformulating time series forecasting, such as electric vehicles charging demand (Qu et al., 2024), taxi and bike usage demand (Liu et al., 2024c), and package delivery demand (Nie et al., 2025a).
@@ 7.4 Summary and outlook
The integration of LLMs as decision facilitators has introduced remarkable capabilities in transportation systems, spanning decision-making, action guidance, and spatial-temporal forecasting. As decision makers, LLMs demonstrate human-like adaptability in traffic signal control, autonomous driving, and route optimization, outperforming traditional methods while offering explainable reasoning. In guiding decisions, LLMs bridge high-level reasoning with low-level control systems through interpretable instructions, enhancing safety and efficiency in complex scenarios. For spatial-temporal forecasting, LLMs decode intricate mobility patterns by unifying multimodal data representations, achieving competitive performance in traffic flow and trajectory prediction through tokenization and fine-tuning strategies. As LLMs evolve from auxiliary tools to central predictors, their integration with ITS techniques such as digital twin platforms and IoT ecosystems will likely promote predictive intelligence in smart transportation applications. Yet, several critical limitations hinder these advances. First, LLMs incur high latency and resource demands, challenging their use in real-time control loops or embedded ITS hardware without significant model compression or architectural redesign. Second, LLMs struggle with exact numerical reasoning and fine-grained geometric constraints inherent to traffic dynamics, leading to suboptimal or unsafe control suggestions when precise calculations matter.
@@
    Data representation and computation: LLMs’ text-centric training limits their effectiveness in processing numerical and geometric data inherent to transportation systems. Although tokenization methods such as spatial-temporal graph tokenizers (Liu et al., 2024e) and vector captioning (Chen et al., 2024b) show promise, fundamental gaps persist in representing continuous physical spaces and performing precise numerical computations. It remains an open question how to develop a customized data representation method that is suitable for multiscale traffic data (Nie et al., 2024b) and adaptable to LLMs.
@@
    Safety and alignment: Critical applications like autonomous driving require rigorous safety guarantees. Current frameworks address this through simulation sandboxes (Fu et al., 2024a) and safety alignment techniques (Yang et al., 2024c; Xie et al., 2025), but real-world deployment demands formal verification methods and enhanced robustness against adversarial inputs (Liu et al., 2024d). Equipping LLMs with safety evaluation modules is a crucial step before practical implementation.
@@ 8.2 Available resources
A Survey of LLM Surveys: https://github.com/NiuTrans/ABigSurveyOfLLMs 2024 GitHub Compiles 150+ surveys across subfields like alignment, safety, and MLLMs.
@@ 8.3 Computational requirement
Pretraining a foundational LLM in a vertical domain such as transportation can be infeasible due to extremely high resource consumption. However, recent advances in PEFT methods such as LoRA (Hu et al., 2022) and QLoRA (Dettmers et al., 2023) have made customization of LLMs widely accessible, allowing mid-range hardware to handle models previously restricted to enterprise-grade infrastructure.
@@ 9.1 Future opportunities of LLM4TR
    Eliciting latent problem-solving abilities for emergent challenges: After being pre-trained on large-scale corpora, LLMs are possessed with potential abilities as general-purpose task solvers. These abilities might not be explicitly exhibited when LLMs perform some specific traffic problems. Therefore, it is useful to design suitable task instructions or specific ICL strategies to elicit such underexplored abilities (Shin et al., 2020; Wei et al., 2022b). Strategic ability elicitation could unlock these capacities. For example, role-playing or persona adoption strategy instructs the model to adopt the persona of a domain expert (e.g., an experienced traveler (Liu et al., 2024f)). This approach leverages the latent knowledge by framing responses in the voice and reasoning style of an expert, which can yield more domain-appropriate answers. The problem decomposition strategy breaks a complex domain-specific problem into a series of simpler sub-problems. Then the compositional tasks can be solved by meta agents. Finally, using domain-contextualized instruction tuning to guide LLMs to focus on relevant domain aspects and follow a structured approach to the problem.
@@
    Model compression for real-time decision-making: Deploying LLMs on resource-constrained edge devices (e.g., onboard vehicle computers, roadside units) necessitates lightweight yet capable models (Liu and Zhao, 2024). Recent advances in model compression suggests a promising pathway. By dynamically pruning non-critical parameters, e.g., tailoring sparsification to preserve traffic-related knowledge such as route optimization while compressing unrelated linguistic knowledge, models can be significantly reduced in size without undermining task-specific performance (Ma et al., 2023a). Furthermore, hardware-aware quantization methods (Lin et al., 2024a), such as mixed-precision and adaptive quantization schemes, are emerging to exploit the computational strengths of modern GPUs/TPUs. One such advancement is the use of 4-bit quantization techniques for efficient LLMs serving (Kim et al., 2023; Zhao et al., 2024b). In addition to quantization, optimizing the KV cache is also crucial for efficient LLM inference (Ge et al., 2023; Liu et al., 2024i; Shi et al., 2024; Liu et al., 2024b). For example, MiniCache (Liu et al., 2024b) compresses the key-value (KV) cache by merging similar states across adjacent layers of LLMs, significantly reducing the memory footprint. These methods can reduce memory footprint and inference latency in a way that is sensitive to the unique requirements of real-time decision-making in transportation applications, without reliance on cloud APIs.
@@ 9.2 Challenges and concerns
Despite the impressive advances in LLMs and their potential to transform transportation systems, the integration and deployment of LLMs into the safety-critical, ethically sensitive and computationally constrained domains raises several challenges that need careful attention. Below, we discuss these concerns related to model biases, domain limitations, computational requirements, data privacy, and ethical implications in depth. Several potential or emerging methods to mitigate them are also mentioned.
@@
    Bias and hallucinations inherent in LLMs: LLMs are typically trained on massive, heterogeneous datasets collected from the internet. Consequently, they may inherit and even amplify biases present in the training data. In transportation contexts, such biases can lead to erroneous predictions or inappropriate recommendations. In addition, the presence of hallucinations, which can cause LLMs to generate factually incorrect information, poses significant risks. In safety-critical systems like autonomous vehicles or traffic control centers, even minor hallucinations can have severe real-world consequences. Recent techniques such as RAG (Guu et al., 2020) can reduce hallucinations by grounding LLMs in factual and domain-specific data. In addition, fine-tuning on curated transportation datasets, use of bias-aware training objectives (Ferrara, 2023), and post-hoc validation layers can help mitigate embedded biases and improve factuality.
@@
    Ensuring rigor and controllability in LLM-driven solutions: While LLMs offer great potential, their integration into transportation systems requires rigor in performance guarantees. A core challenge lies in reconciling the probabilistic, opaque nature of LLMs with the deterministic nature of traffic rules and reliability required for real-time traffic management. Researchers are exploring certifiable validation frameworks that quantify uncertainty bounds (Lin et al., 2023b), enforce context consistency in model outputs, and rigorously test LLM-based solutions against adversarial scenarios (Kumar et al., 2023). To maintain human supervision, explainability tools like counterfactual reasoning interfaces (Feder et al., 2021) or rule-constrained decoding methods should be mandated, allowing operators to monitor how LLM-derived recommendations align with controllability and operational constraints.
@@
    Domain-specific problem-solving and numeric computing limitations: While LLMs excel at processing and generating natural language, their capabilities in precise numerical computing and domain-specific problem solving remain limited. Transportation problems often include precise and quantitative analysis that depend on accurate numerical computations or mathematical optimization, such as real-time traffic prediction, equilibrium analysis, and dynamic routing. Standard LLMs are not inherently designed for these tasks, which can lead to suboptimal performance when used in isolation. Hybrid models that combine LLMs with symbolic solvers (Pan et al., 2023), constraint programming, or external simulation tools (Gao et al., 2024a) are emerging. Alternatively, tool-using LLMs (Nakano et al., 2021; Schick et al., 2023) that call external APIs or plug-ins (e.g., for solving equations or optimization) show promise for extending their capabilities beyond language.
@@
    Computational demand for adapting LLMs: The impressive performance of LLMs comes at a high computational cost. Even fine-tuning and adapting these models on user-defined datasets necessitates prohibitive computational resources. Previous data-driven transportation modeling typically required only moderate or small computational overhead. Thus, computation in the era of LLMs becomes a technical bottleneck for transportation researchers and practitioners with limited hardware computing capacity. PEFT methods such as LoRA, adapter modules, and prompt tuning dramatically reduce resource requirements. Lightweight open-source models or cloud-hosted inference APIs can also enable access to powerful LLMs without full deployment. While API-based solutions offer an accessible approach to using pre-trained LLMs provided by third-part companies, the limited access and model agnosticism largely limits the ways in which capabilities of LLMs can be explored and applied.
@@
    Privacy and security concerns: LLMs require large amounts of data, some of which may be sensitive or personal such as real-time location data, travel patterns, or driver behavior information. This raises significant privacy concerns when LLMs memorize and leak personal information. In addition to privacy issues, there are also security risks related with deploying LLMs as decision makers in critical infrastructure. For instance, cyberattacks using adversarial prompts can manipulate LLM-based traffic controller. Such adversarial vulnerabilities of LLMs are adverse for user’s security and system’s functionality. Differential privacy, federated learning, and on-device inference can help protect user data when using LLMs (Yao et al., 2024b). In terms of security, recent work explores adversarial robustness training and input sanitization to defend against prompt injection attacks in safety-critical deployments (He and Vechev, 2023; Das et al., 2025).
@@
    Ethical, legal, and equitable concerns: Finally, the integration of LLMs in transportation systems raises several ethical, legal, and equitable issues. The black-box nature of many LLMs complicates interpretability, making it difficult to trace decisions or explain errors when they occur. This opacity is problematic in scenarios where transparency is required for public safety and trust, such as autonomous vehicles. Furthermore, there are concerns about equitable access. Biased outputs of LLMs may disproportionately affect underrepresented or vulnerable groups. For example, route optimization models may prioritize routes through lower-income neighborhoods as "acceptable" congestion zones. While the high cost of deploying and maintaining advanced LLMs may also widen the gap between well-funded urban centers and less affluent regions. Lastly, legal liability in the event of system failures remains largely undefined. If an LLM-recommended merging strategy causes a collision of autonomous vehicles, is the responsibility with the model developer, traffic agency, or vehicle original equipment manufacturer? Addressing these challenges will require interdisciplinary collaboration among technologists, policymakers, and ethicist to create guidelines and practices for the responsible use of LLMs in transportation, such as transparent model auditing (Mökander et al., 2024), impact assessments, and equity-aware objective functions during development. On the legal front, ongoing research explores frameworks for assigning liability and setting regulatory standards for LLM-driven decision systems (Hacker et al., 2023).